# How Numerical Precision Affects Mathematical Reasoning Capabilities of LLMs

### v1

@@ Appendix D Proofs for Section 4
For any fixed integers $p$, there exist constant-precision Transformers with constant depth $L$ and hidden dimension $d=O(n^{2})$ that can solve the $\operatorname{ADD}(n,p)$ task.
For ease of reading, we first describe an algorithm to perform $\operatorname{ADD}(n,p)$ (Algorithm 1) and prove the correctness, then construct a Transformer with the given configurations in Theorem 4.1 to simulate Algorithm 1.
Input : Two $p$-adic numbers ${\bm{a}},{\bm{b}}$ with length $n_{1},n_{2}$
Output : The sum of the inputs ${\bm{o}}$, containing $(n+1)$ bits where $n=\max(n_{1},n_{2})$
1 $a_{n}=0,b_{n}=0$;
@@
3       Compute the carry-on bits ${\bm{c}}$:
4       $i_{\wedge}=\max\{j\leq i\mid a_{i}+b_{i}\geq p\}$;
5       $i_{\vee}=\max\{j\leq i\mid a_{i}+b_{i}\leq p-2\}$;
@@ 6       $c_{i}={\bm{1}}_{i_{\wedge}>i_{\vee}}$;
9Compute the outcome ${\bm{o}}$: $o_{i}=(a_{i}+b_{i}+c_{i-1})\operatorname{mod}p$;
Algorithm 1 $\operatorname{ADD}(n,p)$ Algorithm
Lemma D.1 (An algorithm to perform $\operatorname{ADD}(n,p)$).
@@
Given two $n$-bit $p$-adic numbers ${\bm{a}}$ and ${\bm{b}}$, we can compute the carry-over bits ${\bm{c}}=(c_{n},\cdots,c_{1})$ as follows:
@@
By removing the recursive computation, we can compute the carry-over by the following formulas:
$i_{\wedge}=\max\{j\leq i\mid a_{i}+b_{i}\geq p\},$ (4)
$i_{\vee}=\max\{j\leq i\mid a_{i}+b_{i}\leq p-2\},$
@@
Or equivalently,
@@
In Equation 4, $i_{\wedge}$ indicates the closest bit that can contribute an additional carry to the higher bits up to the $i$ th bit, while $i_{\vee}$ identifies the highest bit lower than $i$, such that the carry generated by the bits below $i_{\vee}$ does not affect the bits higher than $i_{\vee}$. Therefore, the carry-over $c_{i}=1$ if and only if $i_{\wedge}<i_{\vee}$.
Then after computing the carryover bits, we can compute the sum of the input integers as follows:
@@
Therefore, the output ${\bm{o}}$ is exactly the sum of the integers and the Algorithm 1 outputs $\operatorname{ADD}({\bm{a}},{\bm{b}})$ for all ${\bm{a}},{\bm{b}}\in\{0,1\}^{n}$. ∎
@@
Now, we demonstrate that a constant-precision transformer, with a constant depth $L$, a fixed number of attention heads, and $O(n^{2})$ hidden dimensions, is capable of simulating Algorithm 1. Consequently, this model can accurately generate correct output for any input integers ${\bm{a}},{\bm{b}}$.
Initial Embeddings: The total length of the input sequence is no longer than $2(n+1)$. We categorize the tokens into two classes: number tokens ($10,1,\cdots,p-$ 1) and auxiliary tokens ($+$, $=$, <SOS> and <EOS>).
  • Embedding of input token $a_{i}$: ${\bm{u}}_{a,i}^{0}=(a_{i}{\bm{e}}_{i+1},{\bm{0}},-1,{\bm{0}},0,1,1)$.
  • Embedding of input token $b_{i}$: ${\bm{u}}_{b,i}^{0}=({\bm{0}},b_{i}{\bm{e}}_{i+1},-1,{\bm{0}},0,2,1)$.
  • Embedding of output token $o_{i}$: ${\bm{u}}_{o,i}^{0}=({\bm{0}},{\bm{0}},o_{i},{\bm{e}}_{i+1},0,3,-1)$.
  • Embedding of “+” token: ${\bm{u}}_{+}^{0}=({\bm{0}},{\bm{0}},-1,{\bm{0}},0,4,-1)$.
  • Embedding of “=” token: ${\bm{u}}_{=}^{0}=({\bm{0}},{\bm{0}},-1,{\bm{0}},1,5,-1)$.
  • Embedding of <SOS> token: ${\bm{u}}_{\texttt{<SOS>}}=({\bm{0}},{\bm{0}},-1,{\bm{0}},0,6,-1)$.
  • Embedding of <EOS> token: ${\bm{u}}_{\texttt{<EOS>}}=({\bm{0}},{\bm{0}},0,{\bm{0}},0,3,-1)$.
where ${\bm{e}}_{i}\in{\mathbb{R}}^{n+1}$ is one-hot vector. Suppose we have $2s$-bit precision and define $B_{s}=2^{s}-2^{-s}$. Recall that, the maximum number we can represent is $B_{s}$, while the minimum is $-B_{s}$.
Block 1. The first block of the Transformer will COPY the value of $a_{i},b_{i}$ to position of outputs. We use the attention head to finish the COPY operation. We set the query, key and value as follows:
  • Query: ${\bm{q}}=B_{s}$.
@@
  • Value: ${\bm{v}}={\bm{u}}^{0}[1,\cdots,2n+2]$, i.e., ${\bm{v}}=(a_{i}{\bm{e}}_{i+1},{\bm{0}})$ for input ${\bm{a}}$, ${\bm{v}}=({\bm{0}},b_{i}{\bm{e}}_{i+1})$ for input ${\bm{b}}$, and ${\bm{v}}={\bm{0}}$ otherwise.
Since we have only constant precision here, we need to check the attention values carefully. The attention value (without normalization) will be $B_{s}$ for tokens $a_{i},b_{i}$, and $-B_{s}$ otherwise. By Lemma C.8, we can get $\exp(B_{s})=B_{s},\exp(-B_{s})=0$, and the normalization term is $2nB_{s}=B_{s}$, indicating the attention weights will be $1$ for token $x_{i},y_{i}$, and 00 otherwise. Thus the attention output at the position of output tokens is always $(a_{0},\cdots,a_{n},b_{0},\cdots,b_{n})$.
Block 2. The second block of the Transformer uses MLPs to calculate the output ${\bm{o}}$ using Algorithm 1. We will finish these calculations following the steps below:
  • Calculate $r_{i}=a_{i}+b_{i}$ for $i=0,\cdots,n$. This can be implemented by a MLP with constant hidden dimension. Here, we require $B_{s}\geq 2p$ to avoid overflow of $r_{i}$.
  • Calculate $f_{i}={\bm{1}}_{r_{i}\geq p}$ and $g_{i}={\bm{1}}_{r_{i}\geq p-2}$. By Lemma C.9, we have
    $f_{i}=\frac{\text{GeLU}[B_{s}\cdot(2r_{i}-2p+1)]}{\text{GeLU}(B_{s})},\;g_{i}=\frac{\text{GeLU}[B_{s}\cdot(2r_{i}-2p+5)]}{\text{GeLU}(B_{s})}.$
@@
  • Calculate $c_{i}$ by the following formula:
@@
    Notice that
    $\bigvee_{1\leq i\leq\gamma}\alpha_{i}=\frac{\text{GeLU}\left[B_{s}\left(\sum_{i=1}^{\gamma}\alpha_{i}\right)\right]}{\text{GeLU}(B_{s})},\;\bigwedge_{1\leq i\leq\gamma}\alpha_{i}=1-\bigvee_{1\leq i\leq\gamma}(1-\alpha_{i}),$
    which implies that $\vee$,$\wedge$ , ∧ can be implemented by constant-depth, constant-precision MLP with constant hidden dimension. Therefore, we can calculate the value of $c_{i}$ using $O(n)$ hidden dimension.
  • Calculate $o_{i}=a_{i}+b_{i}+c_{i-1}$ for $i=0,\cdots,n$. This can be implemented by a MLP with constant hidden dimension. Again, we require $B_{s}\geq 2p$ to avoid overflow of $o_{i}$.
Since we need to calculate $r_{i},f_{i},g_{i},c_{i}$ for any $i$, this requires the hidden dimension of $O(n^{2})$.
Block 3. Finally, this block filters out the token $o_{i}$ from ${\bm{o}}$. Let’s consider the token $o_{i+1}$, where $i\in\{0,\cdots,n-1\}$, we want to predict the next token $o_{i}$. First, we calculate ${\bm{e}}_{i+1}$ by using ${\bm{e}}_{i+2}$ from the positional embedding of ${\bm{m}}_{o,i+1}^{0}$. Then, we will calculate $o_{i}$ by
@@
Since $x=\text{GeLU}(x)-\text{GeLU}(-x)$, we have
$o_{i}=\sum_{j=1}^{n+1}{\bm{e}}_{i+1}[j]{\bm{o}}[j]=\sum_{j=1}^{n+1}[\text{GeLU}({\bm{e}}_{i+1}[j]-B_{s}(2-2{\bm{o}}[j]))-\text{GeLU}(-{\bm{e}}_{i+1}[j]-B_{s}(2-2{\bm{o}}[j]))],$
which implies we can calculate the value of $o_{i}$ using $O(n)$ hidden dimension. The final output from this layer is characterized by the equation:
@@
Predict Next Token. Given the output embeddings of the last Transformer layer ${\bm{e}}_{o,i}^{3}$, and the word embeddings, the Transformer can simply predict the next token by finding the nearest word embeddings.
In this construction, we only require $B_{s}\geq 4p$, which implies constant precision is sufficient. ∎
@@
For any integers $p$ and $L$, and for any polynomial $f$, there exist problem scales $n$ and $k$ such that no constant-precision autoregressive Transformer with $L$ layers and hidden dimension $d<f(n,k)$ can correctly solve the $\operatorname{IterADD}(n,p)$ task.
@@
Suppose there exists integers $p\geq 2,L$ and polynomial $f$, such that for any problem scales $n,k$, there exists a constant-precision autoregressive Transformer with $L$ layers and hidden dimension $d\leq f(n,k)$ that can correctly solve the $\operatorname{IterADD}(k,n,p)$ task.
Consider $\operatorname{Maj}(b_{1},\cdots,b_{k})$, where $b_{i}\in\{0,1\}$. Here we provide a reduction from $\operatorname{Maj}(b_{1},\cdots,b_{k})$ to $\operatorname{IterADD}(2,k^{\prime},p)$ where $k^{\prime}=p^{\lceil\log_{p}k\rceil}\leq pk$. Let $a_{i}=b_{i}(p^{2}-1)$ for $i=1,\cdots,k$, and
$a_{k+1}+\cdots+a_{k^{\prime}}=p^{\lceil\log_{p}k\rceil+1}-(p^{2}-1)\left\lceil\frac{k}{2}\right\rceil.$
This is feasible since
$p^{\lceil\log_{p}k\rceil+1}-(p^{2}-1)\left\lceil\frac{k}{2}\right\rceil\leq(p^{\lceil\log_{p}k\rceil}-k)(p^{2}-1)$
for $p\geq 2$. Thus we can get
$\operatorname{Maj}(b_{1},\cdots,b_{k})=1\iff\sum_{i=1}^{k}b_{i}\geq\left\lceil\frac{k}{2}\right\rceil\iff\sum_{i=1}^{k^{\prime}}a_{i}\geq p^{\lceil\log_{p}k\rceil+1}\iff o_{\lceil\log_{p}k\rceil+1}>0$
Since a bounded-depth, fixed-precision Transformer with polynomial-size generating a single token falls within the complexity class $\mathsf{AC}^{0}$, by the reduction above we can get $\operatorname{Maj}\in\mathsf{AC}^{0}$, which is a contradiction.
@@
For any integers $p$ and $L$, and for any polynomial $f$, there exist problem scales $n$ and $l$ such that no constant-precision autoregressive Transformer with $L$ layers and hidden dimension $d<f(n,l)$ can correctly solve the $\operatorname{MUL}(n,l,p)$ task.
@@
Suppose there exists integers $p\geq 2,L$ and polynomial $f$, such that for any problem scales $n,l$, there exists a constant-precision autoregressive Transformer with $L$ layers and hidden dimension $d\leq f(n,l)$ that can correctly solve the $\operatorname{MUL}(n,l,p)$ task.
Consider $\operatorname{Maj}(c_{1},\cdots,c_{k})$, where $c_{i}\in\{0,1\}$. Here we provide a reduction from $\operatorname{Maj}(c_{1},\cdots,c_{k})$ to $\operatorname{MUL}(n,l,p)$ where
$n=\left(\lceil\log_{p}k\rceil+1\right)\left(p^{\lceil\log_{p}k\rceil}+\left\lfloor\frac{k}{2}\right\rfloor\right)=O(k\log k),\;l=n+\lceil\log_{p}k\rceil=O(k\log k).$
Let
$k^{\prime}=p^{\lceil\log_{p}k\rceil}+\left\lfloor\frac{k}{2}\right\rfloor,\;c_{k+1}=\cdots=c_{k^{\prime}}=1,$
and
${\bm{a}}=c_{1}\underbrace{0\cdots 0}_{\lceil\log_{p}n\rceil}c_{2}\underbrace{0\cdots 0}_{\lceil\log_{p}n\rceil}\cdots c_{k^{\prime}}\underbrace{0\cdots 0}_{\lceil\log_{p}n\rceil},\;{\bm{b}}=1\underbrace{0\cdots 0}_{\lceil\log_{p}n\rceil}1\underbrace{0\cdots 0}_{\lceil\log_{p}n\rceil}\cdots 1\underbrace{0\cdots 0}_{\lceil\log_{p}n\rceil}.$
Under this construction, we can get
$\operatorname{Maj}(c_{1},\cdots,c_{k})=1\iff c_{1}+\cdots+c_{k}\geq\left\lceil\frac{k}{2}\right\rceil\iff c_{1}+\cdots+c_{k^{\prime}}\geq p^{\lceil\log_{p}k\rceil}\iff o_{l-1}>0$
Since a bounded-depth, fixed-precision Transformer with polynomial-size generating a single token falls within the complexity class $\mathsf{AC}^{0}$, by the reduction above we can get $\operatorname{Maj}\in\mathsf{AC}^{0}$, which is a contradiction.

### v2

@@ Appendix D Proofs for Section 4
In this section, we present the formal proofs of the theorems stated in Section 4. Before delving into the proofs, we revisit the role of the tokenizer ${\bm{T}}_{c}$. As established in Section B.4 and Proposition B.6, it suffices to focus on the case of ${\bm{T}}_{1}$, where both the inputs and outputs are tokenized into single digits. This simplification is key to the subsequent analysis and constructions.
D.1 Proof for Theorem 4.1
@@
Fix integers $p\geq 2$ and $c\in\mathbb{N}^{*}$. Consider the tokenizer ${\bm{T}}_{c}$ defined in Eq. 1 for processing the input and output sequences. There exist constant-precision Transformers with constant depth $L$ (independent of $n$) and hidden dimension $d=O(n^{2})$ that can solve the $\operatorname{ADD}_{p}(n)$ task.
To aid readability, we first describe an algorithm to perform the $\operatorname{ADD}_{p}(n)$ task (Algorithm 1) and prove its correctness. Subsequently, we construct a Transformer with the specified configuration in Theorem 4.1 that simulates Algorithm 1.
Input : Two $p$-adic numbers ${\bm{a}},{\bm{b}}$ with lengths $n_{1}$ and $n_{2}$, respectively.
Output : The sum of the inputs, ${\bm{o}}$, represented as a $p$-adic number with $(n+1)$ digits, where $n=\max(n_{1},n_{2})$.
1 Initialize $a_{n}=0$ and $b_{n}=0$;
@@
3       Compute the carry-on bits ${\bm{c}}$;
4       $i_{\wedge}=\max\{j\leq i\mid a_{j}+b_{j}\geq p\}$;
5       $i_{\vee}=\max\{j\leq i\mid a_{j}+b_{j}\leq p-2\}$;
@@ 6       $c_{i}={\bm{1}}_{i_{\wedge}>i_{\vee}}$;
9Compute the output digits ${\bm{o}}$: $o_{i}=(a_{i}+b_{i}+c_{i-1})\operatorname{mod}p$;
Algorithm 1 $\operatorname{ADD}_{p}(n)$ Algorithm
Lemma D.1 (An algorithm to perform $\operatorname{ADD}_{p}(n)$).
@@
Consider two $n$-bit $p$-adic numbers ${\bm{a}}$ and ${\bm{b}}$. The carry-over bits ${\bm{c}}=(c_{n},\cdots,c_{1})$ can be computed recursively as follows:
@@
To avoid the recursive computation, the carry-over bits can be expressed in closed form as:
$i_{\wedge}=\max\{j\leq i\mid a_{j}+b_{j}\geq p\},$ (5)
$i_{\vee}=\max\{j\leq i\mid a_{j}+b_{j}\leq p-2\},$
@@
Alternatively, the carry-over bits can be expressed equivalently as:
@@
In Equation 5, $i_{\wedge}$ identifies the largest bit index less than or equal to $i$ that contributes a carry to higher bits, while $i_{\vee}$ identifies the largest bit index less than or equal to $i$ such that the carry generated below $i_{\vee}$ does not propagate beyond $i_{\vee}$. Thus, the carry-over bit $c_{i}=1$ if and only if $i_{\wedge}>i_{\vee}$.
After computing the carry-over bits, the sum of the input integers can be computed as:
@@
Therefore, the output ${\bm{o}}$ is exactly the sum of the two input numbers, and Algorithm 1 correctly computes $\operatorname{ADD}_{p}({\bm{a}},{\bm{b}})$ for all ${\bm{a}},{\bm{b}}\in\{0,1\}^{n}$. ∎
@@
We now demonstrate that a constant-precision Transformer, with constant depth $L$, a fixed number of attention heads, and a hidden dimension of size $O(n^{2})$, is capable of simulating Algorithm 1. Consequently, this model can accurately produce the correct output for any pair of input integers ${\bm{a}}$ and ${\bm{b}}$.
Initial Embeddings: The total length of the input sequence is at most $2(n+1)$. We categorize the tokens into two distinct classes: (1) number tokens representing digits ($10,1,\cdots,p-$ 1), and (2) auxiliary tokens for operations and control flow (“$+$”, “$=$”, <SOS>, and <EOS>).
The embeddings for each token are initialized as follows:
  • Embedding of input token $a_{i}$:
    ${\bm{u}}_{a,i}^{0}=(a_{i}{\bm{e}}_{i+1},{\bm{0}},-1,{\bm{0}},0,1,1).$
  • Embedding of input token $b_{i}$:
    ${\bm{u}}_{b,i}^{0}=({\bm{0}},b_{i}{\bm{e}}_{i+1},-1,{\bm{0}},0,2,1).$
  • Embedding of output token $o_{i}$:
    ${\bm{u}}_{o,i}^{0}=({\bm{0}},{\bm{0}},o_{i},{\bm{e}}_{i+1},0,3,-1).$
  • Embedding of the “$+$” token:
    ${\bm{u}}_{+}^{0}=({\bm{0}},{\bm{0}},-1,{\bm{0}},0,4,-1).$
  • Embedding of the “$=$” token:
    ${\bm{u}}_{=}^{0}=({\bm{0}},{\bm{0}},-1,{\bm{0}},1,5,-1).$
  • Embedding of the <SOS> token:
    ${\bm{u}}_{\texttt{<SOS>}}^{0}=({\bm{0}},{\bm{0}},-1,{\bm{0}},0,6,-1).$
  • Embedding of the <EOS> token:
    ${\bm{u}}_{\texttt{<EOS>}}^{0}=({\bm{0}},{\bm{0}},0,{\bm{0}},0,3,-1).$
In each of these embeddings:
  • ${\bm{e}}_{i}\in\mathbb{R}^{n+1}$ is a one-hot vector representing the positional encoding of the token (e.g., digit $a_{i}$) in the sequence.
  • ${\bm{0}}$ is a vector of zeros of appropriate dimensions.
Block 1. The first block of the Transformer performs the COPY operation, which copies the values of $a_{i},b_{i}$ to the positions of the output tokens. This is achieved using the attention mechanism. The query, key, and value are set as follows:
  • Query: ${\bm{q}}=B_{s}$
@@
  • Value: ${\bm{v}}={\bm{u}}^{0}[1,\cdots,2n+2]$, i.e.,
    ${\bm{v}}=\begin{cases}(a_{i}{\bm{e}}_{i+1},{\bm{0}})&\text{for input ${\bm{a}}$,}\\ ({\bm{0}},b_{i}{\bm{e}}_{i+1})&\text{for input ${\bm{b}}$,}\\ {\bm{0}}&\text{otherwise.}\end{cases}$
Since we operate under constant precision, we carefully analyze the attention values. The attention value (before normalization) is $B_{s}$ for tokens $a_{i},b_{i}$ and $-B_{s}$ otherwise. By Lemma C.8, we know $\exp(B_{s})=B_{s}$ and $\exp(-B_{s})=0$. The normalization term for attention is $2nB_{s}=B_{s}$, so the attention weights are $1$ for tokens $a_{i},b_{i}$ and 00 otherwise. As a result, the attention output at the positions of the output tokens is always $(a_{0},\cdots,a_{n},b_{0},\cdots,b_{n})$.
Block 2. The second block of the Transformer uses MLPs to compute the output ${\bm{o}}$ based on Algorithm 1. The calculations proceed in the following steps:
  • Compute $r_{i}=a_{i}+b_{i}$ for $i=0,\cdots,n$.
    This can be implemented using an MLP with constant hidden dimension. To avoid overflow of $r_{i}$, we require $B_{s}\geq 2p$.
  • Compute $f_{i}={\bm{1}}_{r_{i}\geq p}$ and $g_{i}={\bm{1}}_{r_{i}\geq p-2}$.
    Using Lemma C.9, we can calculate:
    $f_{i}=\frac{\text{GeLU}[B_{s}\cdot(2r_{i}-2p+1)]}{\text{GeLU}(B_{s})},\quad g_{i}=\frac{\text{GeLU}[B_{s}\cdot(2r_{i}-2p+5)]}{\text{GeLU}(B_{s})}.$
@@
  • Compute $c_{i}$ using the formula:
@@
    Notice that:
    $\bigvee_{1\leq i\leq\gamma}\alpha_{i}=\frac{\text{GeLU}\left[B_{s}\cdot\left(\sum_{i=1}^{\gamma}\alpha_{i}\right)\right]}{\text{GeLU}(B_{s})},\quad\bigwedge_{1\leq i\leq\gamma}\alpha_{i}=1-\bigvee_{1\leq i\leq\gamma}(1-\alpha_{i}).$
    These formulas imply that the $\vee$ and $\wedge$ operations can be implemented using a constant-depth, constant-precision MLP with constant hidden dimension. Therefore, $c_{i}$ can be computed using $O(n)$ hidden dimension.
  • Compute $o_{i}=a_{i}+b_{i}+c_{i-1}$ for $i=0,\cdots,n$.
    This computation can also be implemented with a constant hidden dimension. Again, we require $B_{s}\geq 2p$ to avoid overflow of $o_{i}$.
Since we need to compute $r_{i},f_{i},g_{i},c_{i}$ for all $i$, the hidden dimension of this block is $O(n^{2})$.
Block 3. This block filters out the token $o_{i}$ from ${\bm{o}}$. Specifically, for the token $o_{i+1}$, where $i\in\{0,\cdots,n-1\}$, we predict the next token $o_{i}$.
First, we calculate the positional embedding ${\bm{e}}_{i+1}$ using ${\bm{e}}_{i+2}$ from the positional embedding of ${\bm{m}}_{o,i+1}^{0}$. Then, we compute $o_{i}$ as:
@@
Using the property $x=\text{GeLU}(x)-\text{GeLU}(-x)$, this can be expanded as:
$o_{i}=\sum_{j=1}^{n+1}{\bm{e}}_{i+1}[j]{\bm{o}}[j]=\sum_{j=1}^{n+1}\left[\text{GeLU}({\bm{e}}_{i+1}[j]-B_{s}(2-2{\bm{o}}[j]))-\text{GeLU}(-{\bm{e}}_{i+1}[j]-B_{s}(2-2{\bm{o}}[j]))\right].$
Thus, $o_{i}$ can be calculated using $O(n)$ hidden dimension. The final output from this layer is given by:
@@
Predict Next Token. Given the output embeddings from the last Transformer layer, ${\bm{e}}_{o,i}^{3}$, and the word embeddings, the Transformer predicts the next token by finding the nearest word embedding.
Precision Requirements. In this construction, we require $B_{s}\geq 4p$, which guarantees that constant precision is sufficient for all computations. ∎
D.2 Proof for Theorem 4.2
@@
Fix integers $p\geq 2$ and $c,L\in\mathbb{N}^{*}$. Consider the tokenizer ${\bm{T}}_{c}$ defined in Eq. 1 for processing the input and output sequences. For any polynomial $f$, there exist problem scales $n$ and $k$ such that no constant-precision autoregressive Transformer with $L$ layers and hidden dimension $d<f(n,k)$ can correctly solve the $\operatorname{IterADD}_{p}(n,k)$ task.
@@
Assume, for the sake of contradiction, that there exist integers $p\geq 2$, $L$, and a polynomial $f$, such that for all problem scales $n$ and $k$, there exists a constant-precision autoregressive Transformer with $L$ layers and hidden dimension $d\leq f(n,k)$ that can solve the $\operatorname{IterADD}_{p}(n,k)$ task correctly.
We now consider the majority function $\operatorname{Maj}(b_{1},\dots,b_{k})$, where $b_{i}\in\{0,1\}$. To establish the contradiction, we construct a reduction from $\operatorname{Maj}(b_{1},\dots,b_{k})$ to $\operatorname{IterADD}_{p}(2,k^{\prime})$, where $k^{\prime}=p^{\lceil\log_{p}k\rceil}\leq pk$. Specifically, let $a_{i}=b_{i}(p^{2}-1)$ for $i=1,\dots,k$, and define the remaining terms as follows:
$a_{k+1}+\dots+a_{k^{\prime}}=p^{\lceil\log_{p}k\rceil+1}-(p^{2}-1)\left\lceil\frac{k}{2}\right\rceil.$
This construction is feasible because:
$p^{\lceil\log_{p}k\rceil+1}-(p^{2}-1)\left\lceil\frac{k}{2}\right\rceil\leq(p^{\lceil\log_{p}k\rceil}-k)(p^{2}-1),$
which holds for $p\geq 2$. Consequently, the following equivalence relationships hold:
$\operatorname{Maj}(b_{1},\dots,b_{k})=1\iff\sum_{i=1}^{k}b_{i}\geq\left\lceil\frac{k}{2}\right\rceil\iff\sum_{i=1}^{k^{\prime}}a_{i}\geq p^{\lceil\log_{p}k\rceil+1}\iff o_{\lceil\log_{p}k\rceil+1}>0,$
where $o_{\lceil\log_{p}k\rceil+1}$ is the output token corresponding to the final layer of the Transformer.
Now, observe that a bounded-depth, fixed-precision decoder-only Transformer with polynomial hidden dimension, which generates a single token, operates within the complexity class $\mathsf{AC}^{0}$. However, by the reduction above, solving $\operatorname{IterADD}_{p}(2,k^{\prime})$ implies the ability to compute $\operatorname{Maj}$. This leads to a contradiction, as $\operatorname{Maj}\notin\mathsf{AC}^{0}$.
Thus, no constant-precision autoregressive Transformer with $L$ layers and hidden dimension $d\leq f(n,k)$ can solve the $\operatorname{IterADD}_{p}(n,k)$ task in general. ∎
D.3 Proof for Theorem 4.3
@@
Fix integers $p\geq 2$ and $c,L\in\mathbb{N}^{*}$. Consider the tokenizer ${\bm{T}}_{c}$ defined in Eq. 1 for processing the input and output sequences. For any polynomial $f$, there exist problem scales $n$ and $l$ such that no constant-precision autoregressive Transformer with $L$ layers and hidden dimension $d<f(n,l)$ can correctly solve the $\operatorname{MUL}_{p}(n,l)$ task.
@@
Assume, for contradiction, that there exist integers $p\geq 2$, $L$, and a polynomial $f$, such that for all problem scales $n$ and $l$, there exists a constant-precision autoregressive Transformer with $L$ layers and hidden dimension $d\leq f(n,l)$ that can correctly solve the $\operatorname{MUL}_{p}(n,l)$ task.
Now, consider the majority function $\operatorname{Maj}(c_{1},\dots,c_{k})$, where $c_{i}\in\{0,1\}$. We construct a reduction from $\operatorname{Maj}(c_{1},\dots,c_{k})$ to $\operatorname{MUL}_{p}(n,l)$. Specifically, let
$n=\left(\lceil\log_{p}k\rceil+1\right)\left(p^{\lceil\log_{p}k\rceil}+\left\lfloor\frac{k}{2}\right\rfloor\right)=O(k\log k),\quad l=n+\lceil\log_{p}k\rceil=O(k\log k).$
We extend $c_{i}$ by defining
$k^{\prime}=p^{\lceil\log_{p}k\rceil}+\left\lfloor\frac{k}{2}\right\rfloor,\quad c_{k+1}=\cdots=c_{k^{\prime}}=1,$
and construct the sequences ${\bm{a}}$ and ${\bm{b}}$ as follows:
${\bm{a}}=c_{1}\underbrace{0\cdots 0}_{\lceil\log_{p}n\rceil}c_{2}\underbrace{0\cdots 0}_{\lceil\log_{p}n\rceil}\cdots c_{k^{\prime}}\underbrace{0\cdots 0}_{\lceil\log_{p}n\rceil},\quad{\bm{b}}=1\underbrace{0\cdots 0}_{\lceil\log_{p}n\rceil}1\underbrace{0\cdots 0}_{\lceil\log_{p}n\rceil}\cdots 1\underbrace{0\cdots 0}_{\lceil\log_{p}n\rceil}.$
Under this construction, the following equivalences hold:
$\operatorname{Maj}(c_{1},\dots,c_{k})=1\iff c_{1}+\cdots+c_{k}\geq\left\lceil\frac{k}{2}\right\rceil\iff c_{1}+\cdots+c_{k^{\prime}}\geq p^{\lceil\log_{p}k\rceil}\iff o_{l-1}>0,$
where $o_{l-1}$ denotes the first output token.
Now, observe that a bounded-depth, fixed-precision decoder-only Transformer with polynomial hidden dimension, which generates a single token, operates within the complexity class $\mathsf{AC}^{0}$. However, by the reduction above, solving $\operatorname{MUL}_{p}(n,l)$ implies the ability to compute $\operatorname{Maj}$. This leads to a contradiction, as $\operatorname{Maj}\notin\mathsf{AC}^{0}$.
Hence, no constant-precision autoregressive Transformer with $L$ layers and hidden dimension $d\leq f(n,l)$ can correctly solve $\operatorname{MUL}_{p}(n,l)$ task for any problem scale $n$ and $l$. ∎